OpenStudy (cheesecakekitten):

PLEASE HELP ME I CANNOT FAIL 9TH GRADE

OpenStudy (cheesecakekitten):

Andy and Sam are saving money to go on their senior trip. The amount of money that Andy will have at the end of each week, w, can be expressed at A(w) = 20w + 300. The amount of money that Sam will have at the end of each week, w, can be expressed at S(w) = 15w + 400. They have decided to combine their savings accounts. Write a function that expresses the total amount, T(w) they have in their savings account at the end of each week. A) T(w) = 5w - 100 B) T(w) = 35w + 700 C) T(w) = -5w + 100 D) T(w) = 300w^2 + 12500w + 120000

OpenStudy (johnnewman):

i think may be d i am not sure

OpenStudy (dumbcow):

If Andy has 100 and Sam has 200 , what is the total? and how did you come up with it?

OpenStudy (lpruitt4):

give me a medal

OpenStudy (cheesecakekitten):

300

OpenStudy (lpruitt4):

yup 300

OpenStudy (cheesecakekitten):

@dumbcow Simplify. 3n2 + 8n ------------ n A) 3n + 8n B) 3n + 8 C) 3n2 + 8n2 D) 3n2 + 8

OpenStudy (dumbcow):

yes its 300 because you "ADD" them together ok for this one, factor out a "n" on top, then cancel it with "n" on bottom

OpenStudy (cheesecakekitten):

wut

OpenStudy (cheesecakekitten):

factor? cancel?

OpenStudy (cheesecakekitten):

9th grade sucks

OpenStudy (dumbcow):

to factor out means to divide by n ... or take one away 3n^2 has 2 n's .... take one away, leaving 1 n 8n has 1 n ..... take one away , leaving 0 n's

OpenStudy (dumbcow):